# Towards Agentic OS: An LLM Agent Framework for Linux Schedulers

### v1

@@ 3.1. The Semantic Gap Problem
Domain Knowledge Gap between developer and user: In cloud environments, DevOps engineers configuring Kubernetes lack insight into workload characteristics (latency-sensitive vs. throughput-oriented), resulting in conservative scheduling. Edge and personal devices face worse challenges. Gamers, creative professionals, and office workers lack kernel expertise for optimization. LLMs can bridge this gap by understanding high-level workload patterns from source code and deployment artifacts, and translating them into concrete scheduling policies.
Technical Complexity of Scheduler Development: Linux scheduler development requires mastering kernel programming with lock-free structures, eBPF verification constraints, and CPU/NUMA architectures. This steep learning curve limits innovation to few kernel experts. LLMs with pre-trained domain knowledge enable rapid and automated scheduler development without human expertise.
Dynamic Workload Adaptation: Modern workloads exhibit complex phase behavior: ML training alternates between compute-intensive forward propagation and communication-heavy gradient synchronization; web traffic varies by orders of magnitude daily; build system parallelism changes with dependencies. Manual reconfiguration and optimization cannot keep pace, while AI agents can adapt policies in real-time, operating 24/7 without breaks, much more availability than human experts.
3.2. Economic Viability of AI-Driven Scheduler Optimization
Scheduler optimization becomes economically viable when generation costs are lower than CPU time savings. For example, with \$0.45 AI generation LLM API cost and 20% performance improvement, a 10-hour workload on a \$2/hour instance already breaks even (\$0.45 cost vs \$4 savings). This fundamentally changes scheduler economics: traditionally, high manual engineering costs meant custom schedulers were only justified for large-scale cloud workloads running on hundreds of machines for months, but AI assistance now makes it economical to optimize even short-lived workloads like CI/CD pipelines, batch jobs, or individual builds running for just hours or days on single machines, democratizing performance optimization for previously uneconomical use cases.
@@ 3.4. Challenges in Applying LLM Agents to Schedulers
Our experiments reveal critical challenges for AI-driven scheduler optimization, especially when fully automated: Performance: How do we ensure AI-generated or configured schedulers outperform existing ones rather than degrading performance? Safety: How do we prevent kernel crashes, soft-lockups, stalls, or starvation while maintaining stability? How can we ensure only minimal privilege needed when development and deployment? Efficiency: The 33-minute generation time and the \$6 cost must drop for practical deployment.
@@ 4.1. Experimental Setup
We evaluate SchedCP on two machines, machine 1 is an 86-core, 172 threads Intel Xeon 6787P with 758GB RAM, NVMe SSDs, 10Gbps network, with 2x 256 GB CXL (Compute Express Link) memory device, 3 numa node, running Linux 6.14 with sched_ext. Machine 2 is an 8-core, 8 threads Intel Core Ultra 7 258V with 30GB RAM, NVMe SSDs, 1 NUMA node, running Linux 6.13 with sched_ext. We test Claude Code (Opus 4) as AI agents to validate framework generality. For each case, we test 3 times and get the average results. In all the experiments, the Agent successfully creates working custom scheduler configurations or eBPF programs.
4.2. Performance Impact of AI-Driven Optimization
We evaluate the SchedCP and sched-agent’s ability of effectively configure existing scheduler and learn after first attempt. The Linux kernel build benchmark compile the kernel 6.14 with tinyconfig and “make -j 172” on machine 1. Figure 1 shows performance improvements across three stages: baseline EEVDF, LLM-configured schedulers, and Iteration-improved configurations. We also compare with basic RL algorithms that have been proposed for scheduler optimization (Corbet, 2025), which only tunes scheduler parameter. The workload shows 1.63x speedup from 13.57s to 8.31s using scx_rusty as the first attemp. After 3 iteration of observe-optimization process, the sched-agent selects the scx_layered scheduler and adds 16% additional gain beyond LLM configuration, with total improvements of 1.79x over baseline EEVDF. In contrast, basic RL approaches show no improvement in our tests, likely because they require hardware-specific retraining, which is costly and time-consuming.
@@ 4.2. Performance Impact of AI-Driven Optimization
4.3. Scheduler Synthesis for Batch Workloads
Figure 3 evaluates the SchedCP and sched-agent ability of generating new and better scheduler from scratch on 8 diverse batch workloads (e.g. file compression, video transcoding, software testing, and data analytics tasks) running on machine 2. To simulate a long-tail distribution, each workload comprised 40 parallel tasks: 39 short and one significantly longer, each as a python script or C/C++ program. The agent consistently identified this pattern, implemented a Longest Job First (LJF) scheduling policy, and achieved an average 20% reduction in end-to-end processing time. The cost for this analysis averaged \$0.15 per workload, based on Claude Opus 4 pricing from August 2025. We note that the powerful Claude Opus agent successfully classified all 8 workloads, whereas the smaller Claude Sonnet model could not. In addition to performance gains, our framework’s optimizations reduced generation costs per iteration: time fell from 33 to 2.5 minutes (a 13x reduction), and the monetary cost dropped from \$6 to \$0.5.

### v2

@@ 3.1. The Semantic Gap Problem
Linux scheduler optimization faces three fundamental barriers. First, a domain knowledge gap exists between developers and users: DevOps engineers lack insight into workload characteristics (latency-sensitive vs. throughput-oriented), while edge/personal device users lack both kernel optimization expertise and understanding of application-specific performance targets. Second, technical complexity of scheduler development requires mastering kernel programming with lock-free structures, eBPF verification constraints, and CPU/NUMA architectures, limiting innovation to few experts. Third, dynamic workload behavior presents complex challenges: ML training alternates between compute-intensive forward propagation and communication-heavy gradient synchronization, web traffic varies by orders of magnitude daily, and build system parallelism changes with dependencies.
Prior RL-based schedulers (Mao et al., 2019a; Qiu et al., 2020; Zhang et al., 2024; Mao et al., 2019b) require extensive training per workload type, lack semantic understanding to transfer knowledge across workloads, and cannot generate new scheduling code. LLMs uniquely bridge these gaps by: (1) understanding natural language requirements and source code semantics without task-specific training, (2) synthesizing correct eBPF schedulers based on automatic workload characterization, (3) reasoning about performance trade-offs and system constraints, and critically, (4) operating in the control plane to generate optimized code that runs natively with negligible runtime overhead, unlike traditional ML models that would cause unacceptable inference latency in the scheduler hot path. This control plane separation represents a key architectural insight: LLMs generate and optimize scheduling policies offline, producing native eBPF code that executes without any ML inference overhead during actual scheduling decisions.
@@ 3.3. Challenges in Applying LLM Agents to Schedulers
Our experiments reveal critical challenges for AI-driven scheduler optimization, especially when fully automated: Performance: How do we ensure AI-generated or configured schedulers outperform existing ones rather than degrading performance? Safety: How do we prevent kernel crashes, soft-lockups, stalls, or starvation while maintaining stability? How can we ensure only minimal privilege needed when development and deployment? This presents a fundamental programming language challenge: synthesizing domain-specific code that must satisfy both general safety properties (memory safety, termination) and domain-specific invariants (fairness, liveness)—a problem that requires sophisticated verification techniques beyond standard compiler checks. Efficiency: The 33-minute generation time and the \$6 cost must drop for practical deployment.
@@ 4.1. Experimental Setup
We evaluate SchedCP on two machines, machine 1 is an 86-core, 172 threads Intel Xeon 6787P with 758GB RAM, NVMe SSDs, 10Gbps network, with 2x 256 GB CXL (Compute Express Link) memory device, 3 numa node, running Linux 6.14 with sched_ext. Machine 2 is an 8-core, 8 threads Intel Core Ultra 7 258V with 30GB RAM, NVMe SSDs, 1 NUMA node, running Linux 6.13 with sched_ext. We test Claude Code (Opus 4) as AI agents to validate framework generality. For each case, we test 3 times and get the average results. To mitigate cache warming effects, we clear the page cache (via sync; echo 3 > /proc/sys/vm/drop_caches) before each run and perform a warm-up run that is excluded from measurements. In all the experiments, the Agent successfully creates working custom scheduler configurations or generates new eBPF programs.
4.2. AI configured schedulers for kernel build and schbench
We evaluate the SchedCP and sched-agent’s ability to both select/configure existing schedulers and generate new ones, as well as learn after the first attempt. The iterative refinement simulates realistic deployment scenarios where schedulers are evaluated over multiple runs before production deployment. Note that the attempt counts measure the iteration of the observe-optimization process in the AI Agents, it does not require the workload to be re-run. The Linux kernel build benchmark compiles the kernel 6.14 with tinyconfig and “make -j 172” on machine 1. Figure 1 shows performance improvements across three stages: baseline EEVDF, initial AI-selected schedulers, and iteratively-refined configurations. We also compare with pre-trained RL algorithms that have been proposed for scheduler optimization (Corbet, 2025) out of the box, which only tunes scheduler parameter. The workload shows 1.63x speedup from 13.57s to 8.31s using scx_rusty as the first attempt. After 3 iterations of observe-optimization process, the sched-agent selects the scx_layered scheduler and adds 16% additional gain beyond LLM configuration, with total improvements of 1.79x over baseline EEVDF. In contrast, basic RL approaches show no improvement in our tests, likely because they require hardware or workload-specific retraining, which is costly and time-consuming.
@@ 4.2. AI configured schedulers for kernel build and schbench
4.3. AI-Generated Schedulers for Batch Workloads
Figure 3 evaluates the SchedCP and sched-agent ability to generate new schedulers from scratch (not merely select existing ones) on 8 diverse batch workloads (e.g. file compression, video transcoding, software testing, and data analytics tasks) running on machine 2. To simulate a long-tail distribution, each workload comprised 40 parallel tasks: 39 short and one significantly longer, each as a python script or C/C++ program. The agent consistently identified this pattern and generated custom eBPF code implementing a Longest Job First (LJF) scheduling policy—a scheduler not present in our repository—achieving an average 20% reduction in end-to-end processing time. The cost for this analysis averaged \$0.15 per workload, based on Claude Opus 4 pricing from August 2025. We note that the powerful Claude Opus agent successfully classified all 8 workloads, whereas the smaller Claude Sonnet model could not. In addition to performance gains, our framework’s optimizations reduced generation costs per iteration: time fell from 33 to 2.5 minutes (a 13x reduction), and the monetary cost dropped from \$6 to \$0.5.